Report from sceduling physicist & STAR problems

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Sun Nov 11 2001 - 09:18:34 EST

  • Next message: Flemming Videbaek: "Re: Shift report 20011111 00:00-08:00"

    As you can see from the attached mail the accesses needed by STAR have been quite necessary, and will diminish the available beam in the coming days.
    
    Please all shifters - read Hiro's e-mail carefully. The procedure is also documented in
    the si-dete description in the procedure book. Please read this carefully.
    
    
    ----- Original Message ----- 
    From: Angelika Drees 
    Sent: Saturday, November 10, 2001 11:27 PM
    Subject: bad news
    
    
      
    Hey everybody, 
    due to some problem with my local PC (with all this beautiful mailing 
    list on it!) I couldn't get in contact with you earlier. I apologize. 
    Let me first start with friday before we get to the real bad news. 
    Realizing that the rates thursday-friday night didn't quite match our 
    expectation we used the end of a store for local and global decoupling 
    at storage. This was followed by unsuccessful octupole studies and a 
    short store (again with poor rates) after which we granted ring access 
    upon experiments request.  While ramping down current busses feeding the 
    STAR magnet sparked and arched causing major damage evaporating  the 
    aluminum welding edges and pieces of the cross bar at the location of 
    the contact.  Parts of the cables nearby melted. In order to bring STAR 
    back into business as quickly as possible significant amount of IR 
    access is needed. This is the tentative schedule: 
    
    * Sunday, 4 h access from 10:00 am -2:00 pm: get the cross bar out of 
       wide angle hall for repair 
    * Tuesday, 18 h, 6:00 am-24:00 pm: get cross bar back in, rewire and 
        check, reassemble, magnet hot test, magnet heat run 
    
    RHIC should run for the other experiments for the rest of the weekend 
    including Veterans Day on monday. Plans for 3m beta* injection on 
    tuesday have to be postponed. Unfortunately today we had one 4 h store 
    only and a STAR access for diagnosing the magnet problem followed by a 
    power dip and tandem failure. At this point we are dealing with RF 
    problems and lost the last ramp at the top due to spontaneous debunching 
    in Blue.  We just recovered from the QLI caused by dumping the debunched 
    beam. 
    
    Polarized protons were accelerated in the booster up to extraction 
    energy into the AGS. No progress was made so far injecting  into the AGS 
    but RF people are currently in and ready to work on it as soon as we 
    have a RHIC store in.  Lets cross fingers, 
      
      
    
                                                       \|/ 
                       Angelika             (.~.) 
                                                  o0(_)0o 
    ----------------------------------------------------------- 
    Angelika Drees (Collider-Accelerator Department, AP group) 
    BNL, Building 911b-210 
    work: (631) 344 2348 
    FAX:  (631) 344 5954 
    home: (631) 474 3884 
      
      
      
    
    -- 
                                                    \|/
                     Angelika                      (.~.)
                                                  o0(_)0o
    -----------------------------------------------------------
    Angelika Drees (Collider-Accelerator Department, AP group)
    BNL, Building 911b-210
    work: (631) 344 2348
    FAX:  (631) 344 5954      
    home: (631) 474 3884
      
    



    This archive was generated by hypermail 2b30 : Sun Nov 11 2001 - 09:14:08 EST